Suppose that the total utility function of a consumer is given by TU(x,y) = 3x2 y and the prices of X and Y are 1 Birr and 2 Birr per unit, respectively. If the income of the consumer is 600 Birr and if he spends all of his income on the consumption of commodities of X and Y, find the optimum amount of X and Y that the consumer will consume at equilibrium and find MRTSx,y.
1
Expert's answer
2021-08-20T09:19:22-0400

Derive the budget constraint:

"I=Px\times X+Py\times Y\\600=X+2\times Y"

The utility maximizing rule is where

"(\frac{MUx}{MUy})=(\frac{Px}{Py})"

"TU(x,y)=3\times x^2\times y"

"MUx=\frac{\partial U}{\partial x}=6\times x\times y"

"MUy=\frac{\partial U}{\partial y}=3\times x^2"

"\frac{Px}{Py}=\frac{1}{2}"

"\frac{6\times x \times y}{3\times x^2}=\frac{1}{2}\\\frac{2\times y}{x}=\frac{1}{2}"

"Y=\frac{x}{4}"

Substitute in the budget constraint:

"600=X+2\times Y"

"600=X+2(\frac{x}{4})"

Multiply both sides by 4:

"2400=4\times X+2\times X\\2400=6\times X\\X=400\\Y=\frac{X}{4}=\frac{400}{4}=100"

"TU(x,y)=(400,100)"

The optimum amount of X and Y that the consumer will consume at equilibrium = 400 and 100

"MRTSxy =\frac{MUx}{MUy}\\MUx=6\times x \times y\\MUy=3\times x^2"

"MRTSxy=\frac{6\times x \times y}{3\times x^2}=\frac{2\times y}{x}=2(\frac{100}{400})=\frac{1}{2}\\MRTSxy=\frac{1}{2}"